MATH 375 – Probability Theory Problem Set Guidelines Fall 2011

Weekly homework assignments will be posted each Friday (or earlier) on the course homepage and collected the following Friday. These assignments will usually consist of 15-25 problems (depending on level of difficulty, how many parts there are to each question, and so forth). Here are some guidelines to follow while doing your assignments, for our mutual benefit:

• Because of the availability of complete solutions to all problem sets after the due date, because every effort will be made to return your graded problem sets in a timely fashion, and for reasons of fairness, no problem sets will be accepted for credit after the announced due date, except in the case of a serious medical situation, family emergency, etc. with prior approval. If you are authorized to hand in a problem set late, I will ask you sign a statement that you have not consulted the reserve solutions in preparing your work.

• Please be neat! Write up the problems in order. Use only on one side of the page (don’t worry about saving paper), and staple your assignment together.

• Leave lots of space for me to write comments.

• Copy out the problem statement before giving the solution. (This should help you make sure you have answered everything in the question.)

• Show enough detail and explanation of what you are doing so that even a confused student in the class could follow your solution. Don’t expect that the reader already knows how to solve the problem. Just writing a numerical answer is never sufficient and will not earn full credit, even if the number is correct.

• Proofs should be written in complete sentences, although your sentences will often contain mathematical symbols and statements. Proofread what you have said to be sure it makes sense.

• No “and then a miracle occurs” moments in proofs, please! If you are left with a gap in an argument that you just can’t bridge, acknowledge it. I will indicate how to fill the gap, if appropriate, in my comments.

• Start the assignment early (i.e. the day it is posted). This will allow you ample time to consult with me or your study group (if you work in one) if you get stuck on some problems. If you start the assignment on the day before it is due, you may be very unhappy!

• Work on making your arguments clear and concise. Make good use of notation and diagrams.

• If you work with other students to solve a problem, you must write up your final solutions independently. Plan on taking scrap notes when you work out the problem first, and then write up your solution in a clean form in your own words, by yourself afterwards. Add a note to your solution listing the other people you consulted.